# CAIRe: Cultural Attribution of Images by Retrieval-Augmented Evaluation

### v1

@@ Abstract
As text-to-image models become increasingly prevalent, ensuring their equitable performance across diverse cultural contexts is critical. Efforts to mitigate cross-cultural biases have been hampered by trade-offs, including a loss in performance, factual inaccuracies, or offensive outputs. Despite widespread recognition of these challenges, [^1] For example, in a well-known 2024 incident Google Gemini’s T2I diversification techniques produced offensively ahistorical images [15] an inability to reliably measure these biases has stalled progress. To address this gap, we introduce CAIRe  [^2] Available for download at github.com/siddharthyayavaram/CAIRE, a novel evaluation metric that assesses the degree of cultural relevance of an image, given a user-defined set of labels. Our framework grounds entities and concepts in the image to a knowledge base and uses factual information to give independent graded judgments for each culture label. On a manually curated dataset of culturally salient but rare items built using language models, CAIRe surpasses all baselines by 28% F1 points. Additionally, we construct two datasets for culturally universal concepts, one comprising of T2I generated outputs and another retrieved from naturally-occurring data. CAIRe achieves Pearson’s correlations of 0.56 and 0.66 with human ratings on these sets, based on a 5-point Likert scale of cultural relevance. This demonstrates its strong alignment with human judgment across diverse image sources.
@@ 1 Introduction
Figure 1: We propose CAIRe, a novel evaluation metric that assesses the degree of cultural relevance of an image, given a user-defined set of labels. Unlike existing methods that assume a definition of culture, we let the user specify the proxy of culture such as countries, religions, or cities to assess over as free-text labels.
Current text-to-image (T2I) models, despite their high-quality outputs, have serious issues in cultural representation and sensitivity. Previous research has reported that they produce culturally homogeneous outputs given under-specified prompts [24, 42] and their outputs are disproportionately biased toward Western cultures, failing to depict global diversity [44]. Further, they generate stereotypical, offensive and factually incorrect outputs when asked to be culturally inclusive [4, 8, 21, 52]. The proliferation of AI-generated content, particularly images, is reshaping our digital media ecosystem, with over 15 billion AI-generated images created since 2022.333https://wired.me/culture/ai-image/ Given this widespread adoption, it is imperative to ensure that these models are inclusive and unbiased towards users from diverse backgrounds.
@@
In this paper we introduce Cultural Attribution of Images by Retrieval (CAIRe), a family of evaluation metrics for visual cultural attribution. CAIRe evaluates images over user-defined cultural labels, by grounding them to entities and concepts in a knowledge-base (§3). Our approach takes as input an image and a list of free-text labels representing cultures of interest. It then outputs a score on a five point scale, indicating the relevance of the image to each label in the culture set (Figure 1). The framework operates in two key stages: First, it grounds the image in real-world concepts and objects, leveraging a massively multilingual and multicultural knowledge base. Second, it utilizes a vision-language model (VLM) to estimate cultural relevance in a retrieval-augmented evaluation setup, where all available information about the recognized concepts is leveraged to estimate this score. Importantly, our method is designed as a flexible framework, allowing users to integrate their preferred knowledge bases or VLMs to suit their specific needs and contexts.
All previous works which have attempted to evaluate geographical [18] or cultural diversity [24] of images assign a single country/region culture label to an image with a binary (relevant / not relevant) score. In contrast, CAIRe estimates the degree of relevance across multiple user-defined culture labels. Due to the absence of test sets with such annotations, we first construct a dataset (§5) of rare and culturally significant items, labeled using GPT-4o [36] and verified with Wikipedia. On this test set CAIRe surpasses baselines by 28% F1 points. After validating CAIRe on this dataset, we evaluate its correlation with human judgments of cultural relevance across a broad range of universal concepts. CAIRe is adept at capturing annotators’ judgments of the relevance of these images to their own cultures, with Pearson’s correlations of 0.56 and 0.66 to these human ratings over generated and natural images, respectively, demonstrating its alignment with human judgment across diverse image sources (§6). In summary, we:
  • Formalize the task of visual cultural attribution—providing graded assessments on how relevant an image is to a user-defined set of free-text culture labels.
  • Build two complementary test sets using our formalism to test visual cultural attribution methods. The first covers obscure concepts spanning a variety of culture proxies and is labeled using GPT-4o, [^4] https://huggingface.co/datasets/cmu-lti/caire-specific while the second covers universal concepts and is annotated with human judgment on a five-point scale. [^5] https://huggingface.co/datasets/cmu-lti/caire-universal
  • Introduce the CAIRe framework built using our design principles, which leverages vision-language models, massive KBs, and image-based retrieval to perform visual cultural attribution.
@@ 3 CAIRe Metric Framework
In this section, we detail how CAIRe performs visual cultural attribution using information about similar images retrieved from a knowledge base (KB). CAIRe processes each image in two steps: (a) Visual Entity Linking (VEL), which retrieves KB entities relevant to the image using embedding-based matching [13], and (b) Cultural Relevance Scoring, which uses VLM judgments over a user-specified set of candidate culture labels. An overview of this pipeline is shown in Figure 2 and its details are below.
@@ 3.1 Visual Entity Linking (VEL)
To enable retrieval, we first index images and text from a chosen KB using a V-L encoder (see Step 0 in Figure 2). Note that the framework design allows users to integrate custom KBs and V-L encoders, making CAIRe generalizable across domains. For this work, we use the massively multilingual BabelNet graph [34] as the KB and mSigLIP 666huggingface.co/google/siglip-so400m-patch16-256-i18n [60] as the V-L encoder. BabelNet contains 6 million entities, each associated with images and text descriptions. For each entity, we compute separate mSigLIP embeddings for the linked images and texts.
@@
where $\mathcal{N}(I)$ denotes the set of top-$K$ retrieved KB entities, $E_{KB}$ are embeddings of KB images, and $\text{sim}(\cdot,\cdot)$ represents cosine similarity. We empirically choose $K=20$ in this work.
@@ 3.2 Cultural Relevance Scoring
Once the input image is linked to relevant KB entities, we assess its cultural relevance using large language and vision-language models (LLMs/VLMs), conditioned on the candidate culture labels in $\mathcal{C}$. Specifically, we use textual descriptions from Wikipedia articles corresponding to retrieved entities, combined with the input image (when applicable), to guide the model’s judgment.
@@
In our numerical scoring setup, we prompt the model to produce an integer score between 1 and 5, based on the input Wikipedia text, the image (if available), and a rubric designed to encourage consistent expert-like assessments. An example of the prompt format is shown in Appendix §8.
To ensure reliable and well-structured outputs, we apply constrained decoding, restricting the model’s output space to valid score tokens (1, 2, 3, 4, 5). The final cultural relevance score is computed as:
@@
where $P(k\mid I,T,c,R)$ represents the model’s predicted probability of score $k$ given the input image $I$, Wikipedia text $T$, candidate culture label $c$, and rubric $R$.
@@ 4 Test Set Curation
Given our formulation of visual cultural attribution as requiring both dynamic definitions of culture and graded relevance judgments (§2), it is crucial to evaluate models on a test set that reflects these criteria. However, existing datasets for assessing cultural and geographic diversity focus exclusively on single, binary country-level annotations [24, 18]7), which are insufficient for our task. To address this gap, we construct two complementary test sets—specific and universal—designed to reflect CAIRe’s desiderata.
The specific set consists of real-world images depicting culturally specific, and often obscure, entities. This set enables us to assess whether CAIRe can reliably capture rare and fine-grained cultural elements. To create labels for this set, we use GPT-4o to assign relevance scores across a diverse set of cultural proxies that extend beyond simple geographic categories, directly supporting desideratum 1 from §2.
@@
The specific set contains 68 concepts [20] produced using GPT-4o. We prompt GPT-4o to provide a list of concepts that are rare, yet significant to different proxies of culture, such as regions, religions, festival, ethnicities, philosophies, etc., as laid out in Adilazuarda et al. [1]. This produces a diverse set of entities across many candidate culture label sets, including countries of the world, ethnicities of Africa, states of India, cities of Indonesia, world religions, Native American tribes, and Bronze Age civilizations. Given one proxy, like cities of Indonesia, an image can have multiple gold labels indicating that its relevant to multiple cities of Indonesia. The complete set is provided in Appendix §C.6. Each generated concept and its labels are validated through an affirmative match to an existing Wikipedia article.
Next, we manually collect CC-licensed images from the web for each concept, ensuring that there is no exact match with images in the KB, [^7] The dataset will be open sourced along with the code to avoid test set contamination. Samples from this test set with its respective culture proxies is shown in Figure 3.
@@
The curated specific test set above includes multi-cultural annotations for each concept, but doesn’t allow us to measure the accuracy of CAIRe in estimating the degree of culture relevance, and how closely it matches human judgment. Hence, we curate a list of universal concepts that are common across multiple cultures (like breakfast, weddings, rituals, etc.) and collect both natural and generated images for selected concepts. Using these images, we run a human rating study on Prolific, [^8] https://www.prolific.com/ where we collect cultural relevance ratings for each image on a scale of 1 to 5. Our data curation process is detailed below:
  • Selecting Concepts: We adopt 20 visually depictable and culturally universal concepts [^9] From a set of 298 human universals available here: https://condor.depaul.edu/~mfiddler/hyphen/humunivers.htm provided by Bhatia et al. [3] (full list in Appendix §C.2). This is to ensure that annotators across diverse geographies are equally competent in understanding these concepts and judging their cultural relevance.
  • Selecting Regions: We recruit annotators from China, India, United States, Brazil, Nigeria, Russia, Mexico, Egypt, Germany, and Indonesia, geographically diverse countries that have many available annotators.
  • Generated data (universal-generated): We use Stable Diffusion 3 Medium (SD3) [11] to generate images. We use a simple prompt format: A realistic photo of CONCEPT in COUNTRY, for all combinations of concepts and countries listed above, producing 200 prompts (20 concepts × 10 countries) and 200 images per model.
  • Natural data (universal-natural): We obtain mSigLIP embeddings for Datacomp-1B [12] and retrieve the top-8 images for the same 200 prompts used in universal-generated. We manually select a single high-quality image from the top-8 images, since retrieval may be noisy for certain prompts.
@@
We evaluate 400 images (200 generated from SD3 (universal-generated) and 200 retrieved from DataComp-1B (universal-natural)). Each image receives a cultural relevance rating on a 1-5 Likert scale by 20 annotators each, across the 10 countries mentioned above. The definition of each rating is shown in Appendix C.3 (Table 12). Each image is assigned a cultural relevance score for each country by averaging the Likert scores given by annotators in said country.
@@
We compute Pearson’s correlation coefficient between the absolute cultural relevance scores assigned by human raters and those predicted by our tool for each image. This metric quantifies the degree of linear association between the predicted scores and human ratings, where a high correlation indicates that our tool not only ranks countries accurately but also assigns scores that align closely with human perceptions of cultural relevance.
@@ 5.1 Baseline 1: Vision-Language Encoders
The first baseline uses embeddings of the input image and a set of textual probe prompts using either CLIP ViT-B/32 [40] or mSigLIP [60] to assign a culture label.
@@ 5.2 Baseline 2: Vision-Language Models
We prompt three recent and popular vision-language models (VLMs) to provide a graded score on cultural relevance: Llama-3.2-11B-Vision-Instruct [16], Qwen2.5-VL-7B-Instruct [54], and Pangea-7B-hf [59]. The prompt includes the target culture along with a description of the relevance levels, as defined in Table 12. We encourage the model to follow chain-of-thought reasoning to obtain best results. The full prompt template is provided in Figure 8.
@@ 5.3 Our Metric: CAIRe
We test LLaMa-3.2-11B, Qwen2.5-VL-7B, and Pangea-7B-hf [16, 54, 59] as different LLM/VLM alternatives in the cultural relevance scoring step of CAIRe3.2).
In the experiments with LLMs, we simply provide the text retrieved in the entity linking (VEL) stage (§3.1) as context to the model, but for VLMs, we can provide both the input image and the retrieved entity information. We experiment with two possible ways of augmenting the LLMs and VLMs with additional retrieved context:
@@ 6 Results and Analysis
Model Img. Wiki. Top-20 Prec. F1 $\Delta$${}_{\textsc{CAIRe}(F1)}$
@@
Table 2: Country-wise Pearson’s correlation with human judgment. (N) represents natural data, while (G) represents generated data. The highest and lowest country-wise values are bold and underlined, respectively. A detailed analysis of the results is provided in §6. Models used are LLaMA-3.2-11B-Vision-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-VL-7B-Instruct, and Pangea-7B.
Table 1 presents the Precision & F1 scores for the CAIRe metrics and baselines on the specific test set. Table 2 presents the Pearson correlation coefficients of each model against the human-labeled cultural relevance judgments in the universal set. Our analysis of CAIRe’s performance against the baselines answers six questions:
Does CAIRe correctly attribute specific entities to their respective cultures?
The specific set checks this using particularly difficult real-world example concepts. Table 1 shows that CAIRe consistently outperforms the open-source LM and VLM baselines at this binary cultural relevance classification task. Among all VLMs (Baselines-2), Qwen2.5-VL-7B-Instruct achieves the highest F1-score of 41.0 without additional context, representing the strongest baseline result. Integrating this VLM into CAIRe significantly improves performance (by 28 F1 points!).
How strongly does CAIRe correlate with human opinions?
@@
Does CAIRe’s model of human opinions generalize across diverse cultures?
@@
How effective is CAIRe in visual entity linking, i.e., matching images to the right entities?
The VEL system successfully matches 86% of instances in the specific set, to their exact labeled concepts, as verified through manual evaluation. In instances where an exact match is not achieved, the retrieved concepts often exhibit semantic or cultural similarity to the query image, with only a minimal number of cases being entirely incorrect. We present a few of these examples in §C.1, with most inaccuracies arising from concepts that are absent in the knowledge base.
To estimate an upper bound on how much VEL performance impacts the overall F1 score in the case of the specific set, we incorporate gold-standard Wikipedia pages as context in our cultural relevance scoring, allowing us to quantify retrieval-induced loss (detailed results in §C.5: Table 10).
Which variant of CAIRe performs best?
On the specific set, Qwen2.5-7B-Instruct achieves the highest performance, narrowly outperforming the VL variant despite not using the image. One possible explanation is that when dealing with difficult entities, Qwen-VL may sometimes conflict with retrieved content, leading to errors. In contrast, Qwen2.5-7B-Instruct, which relies solely on retrieved text, avoids this issue and thus achieves slightly better results.
Additionally, models augmented with the full Wikipedia page of the top entity generally outperform those provided only with the titles of the top 20 retrieved entities. While supplying multiple entity titles can help the model form a broader view, it may also introduce noise and conflicting signals, which can hurt performance on fine-grained cultural judgments.
For the universal set, Qwen2.5-VL-7B-Instruct consistently achieves the highest weighted average Pearson correlation across cultures for both natural and generated images, while also exhibiting the lowest variance across countries—indicating clear superiority among the tested VLMs.
Overall, the Qwen family of models delivers the strongest results with CAIRe, showing a 24–27% absolute F1 improvement over native prompting when augmented with retrieved knowledge.
How does performance vary between natural and generated images?

### v2

@@ Abstract
As text-to-image models become increasingly prevalent, ensuring their equitable performance across diverse cultural contexts is critical. Efforts to mitigate cross-cultural biases have been hampered by trade-offs, including a loss in performance, factual inaccuracies, or offensive outputs. Despite widespread recognition of these challenges, [^1] For example, in a 2024 incident Google Gemini’s T2I diversification techniques produced offensive, historically incorrect images [14] an inability to reliably measure these biases has stalled progress. To address this gap, we introduce CAIRe  [^2] https://github.com/siddharthyayavaram/CAIRE, an evaluation metric that assesses the degree of cultural relevance of an image, given a user-defined set of labels. Our framework grounds entities and concepts in the image to a knowledge base and uses factual information to give independent graded judgments for each culture label. On a manually curated dataset of culturally salient but rare items built using language models, CAIRe surpasses all baselines by 22% F1 points. Additionally, we construct two datasets for culturally universal concepts, one comprising of T2I generated outputs and another retrieved from naturally-occurring data. CAIRe achieves Pearson’s correlations of 0.56 and 0.66 with human ratings on these sets, based on a 5-point Likert scale of cultural relevance. This demonstrates its strong alignment with human judgment across diverse image sources.
Figure 1: We propose CAIRe, a novel evaluation metric that assesses the degree of cultural relevance of an image, given a user-defined set of labels. Unlike existing methods that assume a definition of culture, we let the user specify cultures to assess as free-text labels (photo is of a djembe).
@@ 1 Introduction
Current text-to-image (T2I) models, despite their high-quality outputs, have serious issues in cultural representation and sensitivity. Previous research has reported that they produce culturally homogeneous outputs given under-specified prompts [23, 40] and their outputs are disproportionately biased toward Western cultures, failing to depict global diversity [42]. Further, they generate stereotypical, offensive and factually incorrect outputs when asked to be culturally inclusive [4, 8, 20, 48]. The proliferation of AI-generated content, particularly images, is reshaping our digital media ecosystem, with over 15 billion AI-generated images created since 2022.333https://journal.everypixel.com/ai-image-statistics (last accessed: July 29 2025)
Given this widespread adoption, it is imperative to ensure that these models are inclusive and unbiased towards users from diverse backgrounds.
@@
In this paper we introduce Cultural Attribution of Images with Retrieval (CAIRe), a family of evaluation metrics for visual cultural attribution. CAIRe evaluates images over user-defined cultural labels, by grounding them to entities and concepts in a knowledge-base (§3). Our approach takes as input an image and a list of free-text labels representing cultures of interest. It then outputs a score on a five point scale, indicating the relevance of the image to each label in the culture set (Figure 2). The framework operates in two key stages: First, it grounds the image in real-world concepts and objects, leveraging a massively multilingual and multicultural knowledge base. Second, it utilizes a vision-language model (VLM) to estimate cultural relevance in a retrieval-augmented evaluation setup, where all available information about the recognized concepts is leveraged to estimate this score. Importantly, our method is designed as a flexible framework, allowing users to integrate their preferred knowledge bases or VLMs to suit their specific needs and contexts.
All previous works which have attempted to evaluate geographical [17] or cultural diversity [23] of images assign a single country/region culture label to an image with a binary (relevant / not relevant) score. In contrast, CAIRe estimates the degree of relevance across multiple user-defined culture labels. Due to the absence of test sets with such annotations, we first construct a dataset (§4) of rare and culturally significant items, labeled using GPT-4o [34] and verified with Wikipedia. On this test set CAIRe surpasses baselines by 22% F1 points. After validating CAIRe on this dataset, we evaluate its correlation with human judgments of cultural relevance across a broad range of universal concepts.
CAIRe is adept at capturing annotators’ judgments of the relevance of these images to their own cultures, with Pearson’s correlations of 0.56 and 0.66 to these human ratings over generated and natural images, respectively, demonstrating its alignment with human judgment across diverse image sources (§6). We also demonstrate the use of CAIRe in evaluating cultural diversity across a batch of generated images for a given prompt, highlighting CAIRe’s use as a diagnostic tool to reveal T2I models’ biases.
@@ 3 CAIRe Metric Framework
In this section, we detail how CAIRe performs visual cultural attribution using information about similar images retrieved from a knowledge base (KB). CAIRe processes each image in two steps: (a) Visual Entity Linking (VEL), which retrieves KB entities relevant to the image using embedding-based matching [12], and (b) Cultural Relevance Scoring, which uses VLM judgments over a user-specified set of candidate culture labels. An overview of this pipeline is shown in Figure 2 which we detail above.
@@ 3.1 Visual Entity Linking (VEL)
To enable retrieval, we first index images and text from a chosen KB using a V-L encoder (see Step 0 in Figure 2). Note that the framework design allows users to integrate custom KBs and V-L encoders, making CAIRe generalizable across domains. For this work, we use the massively multilingual BabelNet graph [32] as the KB and mSigLIP  [^4] https://huggingface.co/google/siglip-so400m-patch16-256-i18n [55] as the V-L encoder. BabelNet contains 6 million entities, each associated with images and text descriptions. For each entity, we compute separate mSigLIP embeddings for linked images and text.
@@
where $\mathcal{N}(I)$ denotes the set of top-$K$ retrieved KB entities, $E_{KB}$ are embeddings of KB images, and $\text{sim}(\cdot,\cdot)$ represents cosine similarity. We empirically choose $K=20$ in this work as it balances speed and accuracy. Moreover, $K$ is a parameter that can easily be modified in our code.
@@ 3.2 Cultural Relevance Scoring
Once the input image is linked to relevant KB entities, we assess its cultural relevance using vision-language models (VLMs), conditioned on the candidate culture labels in $\mathcal{C}$. Specifically, we use textual descriptions from Wikipedia articles corresponding to retrieved entities, combined with the input image, to guide the model’s judgment.
@@
In our numerical scoring setup, we prompt the model to produce an integer score between 1 and 5, based on the input Wikipedia text, the image, and a rubric designed to encourage consistent expert-like assessments. An example of the prompt format is shown in Appendix §9.
To ensure reliable and well-structured outputs, we apply constrained decoding, restricting the model’s output space to valid score tokens (1, 2, 3, 4, 5). The final score $s$ is computed as:
@@
Using the model’s predicted probability $P(k\mid I,T,c,R)$ of score $k$ given the input image $I$, Wikipedia text $T$, label $c$, and rubric $R$.
@@ 4 Test Set Curation
Existing datasets used to evaluate cultural and geographic diversity rely solely on single, binary country-level labels [23, 17]7). In contrast, §2 highlights the importance of a dynamic notion of culture—one that allows us to independently measure how relevant each image is to each user-defined cultural category. This renders the label format of current test sets incompatible with our task. To bridge this gap, we introduce two complementary test sets – specific and universal, each designed to reflect CAIRe ’s core desiderata.
The specific set consists of real-world images depicting culturally specific, and often obscure, entities. This set enables us to assess whether CAIRe can reliably capture rare and fine-grained cultural elements. To create labels for this set, we use GPT-4o to assign relevance scores across a diverse set of cultural proxies that extend beyond simple geographic categories, directly supporting desideratum 1 from §2 (examples in Figure 3)
@@
Why the two-part test set? The specific set allows us to evaluate CAIRe on highly granular and flexible cultural definitions—such as ethnic groups in Africa or cities in Indonesia—that go beyond coarse national categories. However, these fine-grained labels are difficult to validate with human annotators, as crowdsourcing platforms typically provide only country-level metadata. To address this limitation, we introduce the universal set, which focuses on broad, cross-culturally recognizable concepts (e.g., breakfast, clothing) for which we can reliably obtain human judgments across many regions. Together, the two sets enable us to assess both CAIRe’s ability to understand nuanced cultural definitions and its alignment with human evaluations where such validation is feasible.
@@
The specific set contains 68 concepts [19] produced using GPT-4o. We prompt GPT-4o for rare, culturally distinctive, and visualizable concepts according to a range of cultural proxies [1]. These proxy culture sets include countries of the world, ethnicities of Africa, states of India, cities of Indonesia, world religions, Native American tribes, and Bronze Age civilizations. Within a proxy set, some concepts may be applicable to multiple cultures, for example shakshuka to the cuisines of several Maghrebi and Levantine countries. The complete set is provided in Appendix §C.6. Each generated concept and its labels are validated through an affirmative match to an existing Wikipedia article.
We manually collect online creative commons images for each concept, excluding any images present in our KB index to avoid test set contamination. Sample entities are shown in Figure 3.
@@
While the specific test set allows us to test CAIRe’s performance on real, rare entities, it does not permit evaluating graded relevance judgments. Hence, we curate a universal test set, consisting of 20 visually depictable and culturally universal concepts from [3]. This test set contains both synthetic (universal-generated) and real (universal-retrieved) images of each of these 20 concepts for 10 countries: China, India, the USA, Brazil, Nigeria, Russia, Mexico, Egypt, Germany, and Indonesia.
We generate the synthetic images using Stable Diffusion 3 Medium (SD3) [10] using the prompt A realistic photo of [CONCEPT] in [COUNTRY]. Using mSIGLIP embeddings we gather images from Datacomp-1B according to this same prompt to acquire the natural images in our test set, which we then manually filter for high-quality photos.
We then hire annotators from these same 10 countries using Prolific [^5] https://www.prolific.com/, and collect cultural relevance ratings for each country, for each image on a five-point Likert scale.
@@
We score 400 total images—200 each in universal-generated and universal-retrieved. We attempt to hire 20 annotators from all 10 countries to score each image on the 1-5 Likert scale. This way every image receives relevance scores from annotators in every country, ensuring that we not only assess CAIRe’s ability to measure relevance, but also irrelevance. (Annotator instructions in Appendix C.4, Table 12). We provide further details about the human annotation in §C.9.
@@
We compute Pearson’s correlation coefficient between averaged absolute cultural relevance scores assigned by human raters and those predicted by our tool for each image. This metric quantifies the degree of linear association between the predicted scores and human ratings, where a high correlation indicates that our tool not only ranks countries accurately but also assigns scores that align closely with human perceptions of cultural relevance.
@@ 5.1 Baseline 1: Vision-Language Encoders
The first baseline uses embeddings of the input image and a set of textual probe prompts using either CLIP ViT-B/32 [38] or mSigLIP [55] to assign a culture label. We include a comparison with alternate V-L Encoders in §C.8.
@@ 5.2 Baseline 2: Vision-Language Models
We prompt three recent and popular open-source vision-language models (VLMs) – Llama-3.2-11B-Vision-Instruct [15], Qwen2.5-VL-7B-Instruct [50], and Pangea-7B-hf [54] – as well as one closed-source frontier model, GPT-4o mini, to provide graded scores on cultural relevance. The prompt includes the target culture along with a description of the relevance levels, as defined in Table 12. We encourage the model to follow chain-of-thought reasoning to obtain best results. The full prompt template is provided in Figure 9.
@@ 5.3 Our Metric: CAIRe
We test LLaMa-3.2-11B-Vision-Instruct, Qwen2.5-VL-7B-Instruct, and Pangea-7B-hf [15, 50, 54] as different VLM alternatives in the cultural relevance scoring step of CAIRe3.2).
In our experiments, we provide the input image along with the text retrieved during the entity linking (VEL) stage (§3.1) as context to the model.
We experiment with two possible ways of augmenting the VLMs with additional retrieved context:
@@ 6 Results and Analysis
[…]
@@
Table 2: Country-wise Pearson’s correlation with human judgment. (N) represents natural data, while (G) represents generated data. The highest and lowest model-wise values are bold and underlined, respectively. A detailed analysis of the results is provided in §6. Models used are LLaMA-3.2-11B-Vision-Instruct, Qwen2.5-VL-7B-Instruct, and Pangea-7B-hf.
(a) Gemini Nanobanana Outputs
(b) CAIRe score distribution across 20 countries
Figure 4: CAIRe’s use in evaluating diversity across T2I generated outputs (prompt: a photo of a wedding)
Table 1 presents the Precision & F1 scores for the CAIRe metrics and baselines on the specific test set. Table 2 presents the Pearson correlation coefficients of each model against the human-labeled cultural relevance judgments in the universal set. To analyse results, we ask the following questions:
RQ1: Does CAIRe correctly attribute specific entities to their respective cultures?
The specific set checks this using particularly difficult real-world example concepts. Table 1 shows that CAIRe consistently outperforms both the closed and open-source VLM baselines at this binary cultural relevance classification task.
Among all open-source VLMs (Baselines-2), Qwen2.5-VL-7B-Instruct achieves the highest baseline F1-score of 41.0 without additional context. Integrating this VLM into CAIRe significantly improves performance (by 25 F1 points!). Although GPT-4o mini remains the strongest standalone baseline with an F1-score of 43.1, incorporating CAIRe into weaker open-source models such as LLaMa-3.2-11B-Vision-Instruct and Qwen2.5-VL-7B-Instruct results in performance that surpasses even this top baseline – showcasing the effectiveness of CAIRe.
RQ2: How strongly does CAIRe correlate with human opinions?
@@
RQ3: Does CAIRe’s model of human opinions generalize across diverse cultures?
@@
RQ4: Which variant of CAIRe performs best?
Qwen2.5-VL-7B-Instruct performs best on the specific set, with models using full Wikipedia pages of top entities outperforming those given only the top 20 entity titles. While multiple titles offer broader context, they can introduce noise, harming fine-grained cultural judgments. On the universal set, Qwen2.5-VL-7B-Instruct again leads, showing the highest average Pearson correlation and lowest cross-country variance across both natural and generated images.
Overall, Qwen models perform strongest with CAIRe, achieving a 25% absolute F1 gain in performance on the specific set when enhanced with retrieved knowledge.
RQ5: How does performance vary between natural and generated images?
@@
RQ6: Can CAIRe help identify cultural skews in batches of generated images?
Figure 4 illustrates how CAIRE scores can be aggregated across a set of generated images to reveal broader cultural patterns in model outputs. Rather than evaluating images individually, examining scores at the batch level highlights which cultures a set of generated samples tends to align with more strongly. In this example, several Western countries (e.g., the United States, United Kingdom, Australia, Germany) consistently receive higher scores across multiple images, suggesting the certain cultural perspectives are systematically better represented than others in the model. CAIRe thus provides a structured way to inspect potential cultural skew in model outputs.